How would I find x? My problem is (1/4)^(x-1)=32^(x+3). I've tried plugging in numbers but my teacher showed us a formula I think but I never understood it.
Answer by stanbon(75887)

How would I find x? My problem is (1/4)^(x-1)=32^(x+3). I've tried plugging in numbers but my teacher showed us a formula I think but I never understood it.
------------------
(1/4)^(x-1)=32^(x+3)
---
Change both sides to a common base as follows:
---
[(2^-2)^(x-1)] = [(2^5)^(x+3)]
-----
2^(-2x+2) = 2^(5x+15)
-----
Equate the exponents:
-2x+2 = 5x+15
7x = -13
x = -13/7
===============
Cheers,
Stan H.